OpenStudy (anonymous):

what is the length of the diagonal of a square with area 40cm^2?

OpenStudy (anonymous):

Let x = the side of the square. The diagonal of the square would be \[x \sqrt{2}\] so it would be \[\sqrt{40}\sqrt{2}\] which is \[\sqrt{80} = 4\sqrt{5}\] You can also use the Pytha. Theorem \[\sqrt{\sqrt{40}^2 + \sqrt{40}^2} = \sqrt{80} = 4\sqrt{5}\]

OpenStudy (anonymous):

Yes, what chiaroscuro did.

OpenStudy (anonymous):

pretty confusing

OpenStudy (anonymous):

just stick with the pytha. theorem then do you know how to get the side? if the area is 40 then the square root of that would be one of the sides. so you now you have a isosceles right triangle with to equal sides sqrt(40) just use the pytha theorem, like i showed up there and you get the answer. does it make sense?

OpenStudy (anonymous):

i dont understand how we will get root 80-4root 5

OpenStudy (anonymous):

oh...nw i get it thanx..

OpenStudy (anonymous):

np, sorry, my browser crashed